Please help me solve this word problem. For a set of 3 conseconsecutive odd integers, 3 times the largest integer is 7 less than twice the sum of the other 2 integers. What is the largest integer in the set

Answer by ankor@dixie-net.com(22740)

For a set of 3 consecutive odd integers,
n, (n+2), (n+4)
3 times the largest integer is 7 less than twice the sum of the other 2 integers.
3(n+4) = 2(n + n+2) - 7
3n + 12 = 2(2n + 2) - 7
3n + 12 = 4n + 4 - 7
3n + 12 = 4n - 3
12 + 3 = 4n - 3n
15 = n is the 1st integer
:
What is the largest integer in the set
15 + 4 = 19 is the largest integer
:
:
See if that checks out in the statement
"3 times the largest integer is 7 less than twice the sum of the other 2 integers. "
3(19) = 2(15 + 17) - 7
57 = 2(32) - 7